7x+4=5x+5 type an integer or a fraction
Answer by algebrahouse.com(1659)

7x + 4 = 5x + 5
2x + 4 = 5 {subtracted 5x from both sides}
2x = 1 {subtracted 4 from both sides}
x = 1/2 {divided both sides by 2}

For more help from me, visit: www.algebrahouse.com